If Steven can mix 20 drinks in 5 minutes, Sue can mix 20 drinks in 10 minutes, and Jack can mix 20 drinks in 15 minutes, how much time will it take all 3 of them working together to mix 20 drinks?
Answer by checkley77(12844)

1/5+1/10+1/15=1/x
(6+3+2)/30=1/x
11/30=1/x
11x=30
x=30/11
x=2.73 min. for the three to mix 20 drinks working together.